A pendulum has a period of 0.69 s on Earth. What is its period on Mars, where the acceleration of gravity is about 0.37 that on Earth?
Answer #1

T = 2π√(l/g)

First, find l.

l = (0.69/2π)^2 x 9.8 = 0.12 m

Now, use the value to find the period on Mars

T = 2π √(0.12/9.8 x 0.37) = 1.13 s